A (Ni+2) solution is electrolyzed using a current of 1.25A. What mass of (Ni0 )

plates out in 30 minutes.


1
Expert's answer
2021-09-16T02:01:08-0400

Solution:

Current, 1.25 A

Time, t = 30 min = 1800 s

Quantity of electricity, Q = I × t = (1.25 A) × (1800 s) = 2250 C

Mass of nickel (Ni), m = ??? g

Molar mass of nickel (Ni), M = 58.69 g/mol


Ni2+ + 2e → Nio (2 electrons)


Faraday's 1st law: The mass of a substance produced at an electrode during electrolysis is proportional to the number of moles of electrons (the quantity of electricity) transferred at that electrode.

Faraday's 1st law can be expressed as:

m = (Q × M) / (Z × F)

where:

m = mass of material (g)

M = molar mass of material (g/mol)

Q = quantity of electricity (C)

Z = number of electrons

F = Faraday constant, 96485 (C mol-1)

 

Thus:

m(Ni) = (2250 C × 58.69 g/mol) / (2 × 96485) = 0.684 g

Mass of nickel (Ni) = 0.684 g


Answer: 0.684 g